A balloon contains 14.0 L of air at a pressure of 760 torr. What will be its volume  when the balloon is taken to a depth of 10 ft in a swimming pool ? assume temperature of air and water to be equal  ( density of Hg 13.6 g/ml )



Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

P1V1=P2V2, P1=13600*9.8*0.76,V1=14.0 L, P2=P1+(120*2.54*0.001*1000*9.8), solve for V2

ADIT KUMAR
12 years ago

thanks for your time  

but how did u go about all these calculations ! ?  

could u elaborate ! 

thanx :)